### v1

@@ VIII Conclusion and outlook
with the subleading field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ being non-constant, and the “kinetic” term surviving the infinite $\omega$ limit, per $\omega\left(\nabla\phi/\phi\right)^{2}\simeq\bigl{(}\ \stackrel{{\scriptstyle 1}}{{\varphi}}\>\bigr{)}^{2}$.
The proof presented in Section VII solidifies this behavior, elucidating its origin and demonstrating that it persists irrespective of the EMT trace of the matter source.
The “remnant” field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ emerges as a free massless scalar field, leaving its footprints in the limiting vacuum. In the broader context of Bergmann-Wagoner theory, the possibility of such a “remnant” field emerging and, if so, its role in the Damour-Nordtvedt attractor mechanism would deserve further examination.
@@
For the generic Bergmann-Wagoner action (1), the shift to the Einstein frame, viz. $\sqrt{-\tilde{g}}\left[\tilde{{\cal R}}-\frac{1}{2}\left(\tilde{\nabla}\varphi\right)^{2}...\right]$, entails a Weyl mapping given by
@@
which is exclusive to the gravitational sector. Hence, for a given function $\omega(\phi)$, any properties that are result from the $\sqrt{2\omega(\phi)+3}$ term should be universal, independent of the EMT trace of matter.
We demonstrated the validity of our proof through a specific case—the BD action coupled with a massless Klein Gordon scalar field (the EMT of which has non-vanishing trace). In Sections IV and V, we expanded upon Bronnikov’s approach [26] by utilizing the Einstein frame and the harmonic radial coordinate. Despite the complication of a dilatonic coupling between the Einstein-frame BD field and the Klein-Gordon field, the system of field equations prove to be soluble. We obtained a static spherisymmetric solution for this coupled action, which can be useful for future studies of BD theory sourced by trace-carrying matter. For the purpose at hand, the characteristic signature of $1/\sqrt{\omega}$ is distinctly evident in the final solution.
@@
Outlook #1—Let us stress that the issue tackled in this paper is not just a question of formal mathematics, but that genuine potentially new physical features are at stake, in both the $T=0$ and $T\neq 0$ cases. Indeed, it results from an $\mathcal{O}\left(1/\sqrt{\omega}\right)$ behavior of the BD scalar field that the large $\omega$ Brans-Dicke vacuum spherical solution (i.e. Brans class I if $\omega>-3/2$) does not converge to the GR vacuum spherical solution (i.e. Schwarzschild), but to the GR spherical solution filled by a massless scalar (i.e. the FJNW solution). Especially in the strongly scalarized case, the FJNW solution exhibits drastically specific features with respect to Schwarzschild. Indeed, for a given mass of the field, circular orbits exist for any areal radius (while the areal radius is lower bounded in Schwarzschild), with a diverging orbital frequency measured by a remote observer (while this frequency is upper bounded in Schwarzschild), despite the Einstein gravitational red-shifting [39]. Since the associated gravitational wave (GW) frequency, in the Extreme Mass Ration Inspiral (EMRI) scheme, is twice the orbital one, the potential impact of the effect in the new GW astronomy is obvious. Still in the strongly scalarized case, another qualitative specific features with respect to Schwarzschild concerns the light bending, which experiments a maxima when the impact parameter decreases. This results in the presence of a caustic surface, which yields an infinite amplification of the apparent luminosity of a pointlike star crossing this caustics [40]. Thence, the case of an orbiting star on a properly inclined orbit results in a sequence of luminosity peaks, an effect which does not exist in the Schwarzschild spacetime. In the $T\neq 0$ case, the age of the Universe is not constrained to a fixed value by the present Hubble parameter $H_{0}$ in BD dust filled flat Robertson-Walker cosmology, even in the $\omega\rightarrow 0$ limit, while it is fixed to $2/\left(3H_{0}\right)$ in GR dust filled flat Robertson-Walker cosmology [24]. We plan to visit this issue in the future.
Outlook #2—Whether and/or how the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly manifests in physical situations warrant a separate investigation. In particular, the Robertson parameter $\gamma$ of BD gravity has been derived via the parametrized post-Newtonian (PPN) approximation to be $\gamma_{PPN}=\frac{\omega+1}{\omega+2}$, which lacks the ${\cal O}\left(1/\sqrt{\omega}\right)$ signature. This should prompt a scrutiny of the validity of the PPN result, including the underlying assumptions in its derivation. Notably, a recent study has challenged the weak-field assumption inherent in the PPN approximation [41]. We aim to delve into this matter in a future report [42].
@@ Acknowledgements.
H.K.N. thank Valerio Faraoni for helpfully pointing out a number of crucial references, and Mustapha Azreg-Aïnou, Tao Zhu, Luis A. Correa-Borbonet for stimulating discussions.
—————$\infty$—————–
Appendix A   Exact solution of BDKG action in isotropic coordinate
For many purposes, it can be worth rewriting the exact solution (121)–(123) in isotropic form. Since one is finally interested in the large $\omega$ limit, let us specify to the $\omega>-3/2$ case (which, besides, is the condition for the BD theory to be Ostrogradsky stable [43]).
@@ Appendix A   Exact solution of BDKG action in isotropic coordinate
Let us remark that $\rho=\infty$ for $u=0$. The solution (121)(123) then achieves the form, rescaling $dt$ and $ds$ by a $2K/\sqrt{\chi}$ factor
@@
Appendix B   Exact solution of BDKG action in “standard” coordinate
There exists a convenient expression for the case of $\chi>0$. Let us make a coordinate transform $u(r)$ such that $-\left(\phi g_{tt}\right)\times\left(\phi g_{rr}\right)=1$, viz.
$e^{-2au}\times e^{2au}\frac{\chi^{2}}{\sinh^{4}\sqrt{\chi}u}\left(\frac{du}{dr}\right)^{2}$ $=1$ (169)
From here we obtain:
$\frac{\chi}{\sinh^{2}\sqrt{\chi}u}du=\epsilon\,dr$ (170)
with $\epsilon$ being a sign. Choosing $\epsilon=-$, we get
$r-r_{0}=\frac{\sqrt{\chi}}{\tanh\sqrt{\chi}u}$ (171)
$\Rightarrow u$ $=\frac{1}{\sqrt{\chi}}\text{arctanh}\frac{\sqrt{\chi}}{r-r_{0}}=\frac{1}{2\sqrt{\chi}}\ln\frac{r-r_{0}+\sqrt{\chi}}{r-r_{0}-\sqrt{\chi}}$ (172)
Note: since $\tanh(\sqrt{\chi}\,u)\in(-1,1)$, we must restrict $r$ to be in the range $(-\infty,r_{0}-\sqrt{\chi})\cup(r_{0}+\sqrt{\chi},+\infty)$. Then
$e^{-2au}=\left(\frac{r-r_{0}-\sqrt{\chi}}{r-r_{0}+\sqrt{\chi}}\right)^{\frac{a}{\sqrt{\chi}}}=\left(1-\frac{2\sqrt{\chi}}{r-r_{0}+\sqrt{\chi}}\right)^{\frac{a}{\sqrt{\chi}}}$ (173)
and
$\frac{e^{2au}\chi}{\sinh^{2}\sqrt{\chi}\,u}$ $=\left(1-\frac{2\sqrt{\chi}}{r-r_{0}+\sqrt{\chi}}\right)^{-\frac{a}{\sqrt{\chi}}}\left((r-r_{0})^{2}-\chi\right)$ (174)
Define $r^{\prime}:=r-r_{0}+\sqrt{\chi}$, making $r^{\prime}\in(-\infty,0)\cup(2\sqrt{\chi},+\infty)$, the metric
$ds^{2}$ $=\frac{1}{\phi(r)}\biggl{[}-\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\frac{a}{\sqrt{\chi}}}dt^{2}$
$+\left(1-\frac{2\sqrt{\chi}}{r}\right)^{-\frac{a}{\sqrt{\chi}}}\left(dr^{2}+\left(1-\frac{2\sqrt{\chi}}{r}\right)\,r^{2}\,d\Omega^{2}\right)\biggr{]}$ (175)
The scalar field (in which $\kappa=\frac{\chi-a^{2}}{2\omega+3}\geqslant 0$)
$\phi(r)={\displaystyle\frac{\phi_{0}\,\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\sqrt{\frac{\kappa}{\chi}}}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\sqrt{\frac{\kappa}{\chi}}}\right)^{2}}}$ (176)
The KG field
$\Psi(r)=-\frac{\Psi_{*}\phi_{0}}{2\sqrt{\kappa}}\frac{\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\sqrt{\frac{\kappa}{\chi}}}}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\sqrt{\frac{\kappa}{\chi}}}}\,.$ (177)
If $a=\sqrt{\chi}$, then $\kappa=0$, forcing $\Psi(r)=0$ everywhere. At the same time, the metric is reduced to the Schwarzschild metric
$ds^{2}$ $=-\left(1-\frac{2\sqrt{\chi}}{r}\right)dt^{2}+\left(1-\frac{2\sqrt{\chi}}{r}\right)^{-1}dr^{2}+r^{2}\,d\Omega^{2}$ (178)

### v3

@@ VII Conclusion and outlook
with the sub-leading field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ being non-constant, and the “kinetic” term surviving the infinite $\omega$ limit, per $\omega\left(\nabla\phi/\phi\right)^{2}\simeq\bigl{(}\ \nabla\stackrel{{\scriptstyle 1}}{{\varphi}}\>\bigr{)}^{2}$.
The proof presented in Section IV solidifies this behavior, elucidating its origin and demonstrating that it persists irrespective of the EMT trace of the matter source.
The “remnant” field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ emerges as a free massless scalar field, leaving its footprints in the limiting spacetime configuration as $\omega\rightarrow\infty$. In the broader context of Bergmann-Wagoner theory, the possibility of such a “remnant” field emerging and, if so, its role in the Damour-Nordtvedt attractor mechanism would deserve further examination.
@@
Along the same line of reasoning, for the generic Bergmann-Wagoner action (1), the shift to the Einstein frame, viz. $\sqrt{-\tilde{g}}\left[\tilde{{\cal R}}-\frac{1}{2}\left(\tilde{\nabla}\varphi\right)^{2}...\right]$, entails a Weyl mapping given by
@@
which is also exclusive to the gravitational sector. Hence, for a given function $\omega(\phi)$, any properties that are result from the $\sqrt{2\omega(\phi)+3}$ term should be universal, independent of the EMT trace of matter.
The proof presented in Section IV is general. We further demonstrated its validity through a specific case—the BD action coupled with a massless Klein Gordon scalar field (the EMT of which has non-vanishing trace). In Section III, we expanded upon Bronnikov’s approach [26] by utilizing the Einstein frame and the harmonic radial coordinate. Despite the complication of a dilatonic coupling between the Einstein-frame BD field and the Klein-Gordon field, the system of field equations prove to be soluble. We obtained a static spherically symmetric solution for this coupled action, which can be useful for future studies of BD theory sourced by trace-carrying matter. For the purpose at hand, the characteristic signature of $1/\sqrt{\omega}$ is distinctly evident in the final solution.
@@
Outlook—We discussed the physical implications of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly in Section VI. There are other possibilities that warrant further investigation. In particular, the Robertson parameter $\gamma$ of BD gravity has been derived via the parametrized post-Newtonian (PPN) approximation to be $\gamma_{PPN}=\frac{\omega+1}{\omega+2}$, which lacks the ${\cal O}\left(1/\sqrt{\omega}\right)$ signature. This should prompt a scrutiny of the validity of the PPN result, including the underlying assumptions in its derivation. Notably, a recent study has challenged the weak-field assumption inherent in the PPN approximation [43]. We have also delved into this matter in two recent reports [44, 45].
@@ Acknowledgements.
We are grateful to the anonymous referees for their insightful feedbacks which helped improve the quality of this paper. H.K.N. thank Valerio Faraoni for helpfully pointing out a number of crucial references, and Mustapha Azreg-Aïnou, Tao Zhu, Luis A. Correa-Borbonet for stimulating discussions.
———$\infty$————
Appendix A Field equations of Einstein-Klein-Gordon-dilation (EKGD) action
This Appendix derives the field equations for the EKGD action (38). Upon variation
[…]
For other terms, denoting the symmetric rank-2 tensor $A_{\mu\nu}:=\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$ and $A:=\tilde{g}^{\mu\nu}A_{\mu\nu}$ the trace of $A_{\mu\nu}$, we have
$\delta\left(\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}A_{\mu\nu}\right)$
$=\sqrt{-\tilde{g}}\left(\frac{1}{2}\tilde{g}^{\mu\nu}A-A^{\mu\nu}\right)\delta\tilde{g}_{\mu\nu}+\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\delta A_{\mu\nu}$ (112)
Regarding the tensor $A_{\mu\nu}$:
the first term is
$\delta\left(\partial_{\mu}\theta\partial_{\nu}\theta\right)$ =[~$=\left[\tilde{\nabla}_{\mu}\left(\delta\theta\partial_{\nu}\theta\right)+\tilde{\nabla}_{\nu}\left(\delta\theta\partial_{\mu}\theta\right)\right]-2(\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\theta)\,\delta\theta$ (114)
This results in, up to a total derivative
$\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\delta\left(\partial_{\mu}\theta\partial_{\nu}\theta\right)$ $=-2\sqrt{-\tilde{g}}\tilde{\square}\,\theta\,\delta\theta$ (115)
and
Similarly for Eq. (115), up to a total derivative
$\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}e^{-\theta}\delta\left(\partial_{\mu}\Psi\partial_{\nu}\Psi\right)$
$=2e^{-\theta}\partial_{\mu}\theta\sqrt{-\tilde{g}}\delta\Psi\tilde{\nabla}^{\mu}\Psi-2\sqrt{-\tilde{g}}e^{-\theta}\tilde{\square}\,\Psi\,\delta\Psi$ (117)
leading to
$\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\delta\left(e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi\right)$
$=\sqrt{-\tilde{g}}e^{-\theta}\left[2(\partial_{\mu}\theta\tilde{\nabla}^{\mu}\Psi-\tilde{\square}\,\Psi)\delta\Psi-(\tilde{\nabla}\Psi)^{2}\,\delta\theta\right]$ (118)
Combining all terms, the variation of the action (38) is
[…]
$+\Bigl{[}(2\omega+3)\sqrt{-\tilde{g}}\tilde{\square}\,\theta+\sqrt{-\tilde{g}}e^{-\theta}(\tilde{\nabla}\Psi)^{2}\Bigr{]}\,\delta\theta$ (119)
$+2\sqrt{-\tilde{g}}e^{-\theta}(-\partial_{\mu}\theta\tilde{\nabla}^{\mu}\Psi+\tilde{\square}\,\Psi)\,\delta\Psi$
Imposing the extremality on (119), regarding $\delta\tilde{g}_{\mu\nu}$, we obtain
[…]
Regarding $\delta\theta$, Eq. (119) yields
$(2\omega+3)\,\tilde{\square}\,\theta=-e^{-\theta}(\tilde{\nabla}\Psi)^{2}$ (122)
Regarding $\delta\Psi$, Eq. (119) yields
~$\tilde{\square}\,\Psi=\tilde{\nabla}^{\mu}\Psi\tilde{\nabla}_{\mu}\theta$ (123)
The last two equations can also be neatly cast as
~$\tilde{\nabla}^{\mu}\left(e^{-\theta}\tilde{\nabla}_{\mu}\Psi\right)$ $=e^{-\theta}\left(-\tilde{\nabla}^{\mu}\theta\tilde{\nabla}_{\mu}\Psi+\tilde{\square}\Psi\right)=0$ (125)
When a non-constant Klein-Gordon (KG) field $\Psi$ is present, Eq. (122) requires that the ‘dilaton’ field $\theta$ must also be non-constant.
Appendix B   The new exact solution of BDKG action in isotropic coordinates
For many future purposes, it can be useful to express the exact solution (77)–(79) in isotropic form. Since one is finally interested in the large $\omega$ limit, let us specify to the $\omega>-3/2$ case (which, besides, is the condition for the BD theory to be Ostrogradski stable [46]).
@@ Appendix B   The new exact solution of BDKG action in isotropic coordinates
Let us remark that $\rho=\infty$ for $u=0$. The solution (77)–(79) then achieves the form, rescaling $dt$ and $ds$ by a $2K/\sqrt{\chi}$ factor